When the digits of a two-digit number are reversed, the new number is 9 more than the original number, and the sum of the digits of the original number is 11. What is the original number?

Answer by arunpaul(104)

let the two digit number be xy
such that 10x +y =xy
now as per question
x+y = 11 -----eq1
10y +x = 10x+y+9 ------eq2
or 9y-9x=9
multiply eq1 with9
so we have
9x +9y = 99 ----eq3
add eq3 and eq2 we have
9x +9y = 99
9y -9x = 9
---------------
18y = 108
y = 108/18 = 6
so x = 5 ( by substituting the value of y in eq1)
(Ans)